# Seeing Sound, Hearing Sight: Uncovering Modality Bias and Conflict of AI models in Sound Localization

### v1

@@ 2.2 Experimental conditions
Vision Only (VOnly) exploits multi-modal biases or priors. In this condition, only visual scenes are provided to the AI models, and they must localize the sounding object despite the absence of any meaningful sound. The audio could either be completely silent or filled with random Gaussian noise (mean 0, standard deviation 1). For example, the same visual stimulus as in the congruent condition is presented, but the dog sound is replaced with silence or noise (Fig. 2(e)). We use the same set of 2,920 images as in the congruent condition.
Multi-Instance Localization (MultiInstLoc) follows the same motivation as the cocktail party problem haykin2005cocktail and features several objects of the same category, but the audio corresponds to only one specific instance, testing localization accuracy in a multi-instance scenario. For example, as illustrated in Fig. 2(f), a dog sound is played at the location of the left dog, while both dogs are visually present in the scene. This condition follows the same setup as the Congruent condition, but is more challenging due to the presence of multiple visually relevant objects. We selected images from the test set of the MSCOCO dataset, where the number of object instances within a given image is restricted to between 2 and 5 for the target category. This results in 263,919 image-audio pairs.
@@ 2.3 Human psychophysics experiment
We conducted in-lab psychophysics experiments with 15 participants, collecting a total of 4,050 trials. All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution. Every experiment lasted approximately 40 minutes. The experimental setup is schematically illustrated in Fig. 2(h). Each trial began with a fixation cross displayed for 500 milliseconds, followed by the presentation of an image and audio pair drawn from one of the six experimental conditions (see Sec 2.2). The 6-second audio clip paired with the image stimulus continuously loops until the trial concludes. Participants were instructed to use a computer mouse to click on the perceived location of the sound source within a time limit of 20 seconds, while wearing stereo headphones throughout the experiment. All trials were randomly sampled, and their presentation sequences were shuffled to minimize order effects. If participants failed to respond within the allotted 20 seconds, the trial automatically ended, and the next trial commenced. Instead of relying on the pre-rendered image-audio pairs, we conducted an audio calibration procedure at the start of the experiment to account for individual differences in height and auditory perception. To achieve this, we implemented real-time stereo audio synthesis in Unity. Following calibration, an audio validation task was conducted to ensure successful calibration and adequate spatial hearing accuracy from the participants. For more details about audio calibration and validation, see Supp Sec. S2 and Supp Fig.S4.
Figure 3: Overview of state-of-the-art multi-modal models for sound source localization (SSL). All models take a pair of images and stereo or mono audio as input, processing them separately through vision and audio encoders. For CNN-based models (indicated by blue arrows), feature maps from the final layers of each encoder are extracted, compared via cosine similarity, and used to generate a similarity heatmap. For transformer-based models, output token sequences are obtained from both encoders. Dot products on their token embeddings are calculated, and a heatmap is produced accordingly. During evaluation, SSL accuracy is determined by verifying whether the location of the maximum activation on the heatmaps (red triangles or red tokens) lies within the segmentation mask of the ground truth sounding object (red bounding box).
2.4 Methods and Evaluation Metrics
We benchmarked ten state-of-the-art multi-modal models on the same set of stimuli used in our human experiments. These models include SSLTI ssltie , LVS vggss , FNAC fnac , IS3 is3 , the fine-tuned version of IS3 (IS3-FT), CAVP cavp , AVSegformer gao2024avsegformer , ImageBind imagebind and LanguageBind languagebind . We provide a unified overview of all the SSL models in Fig. 3. We briefly introduce IS3, IS3-FT, and random in the main text. See Supp Sec. S3 for the details about other models.
@@ 2.4 Methods and Evaluation Metrics
IS3 is3 is a dual-stream architecture with 2D Convolutional Neural Networks (CNNs), which processes visual and auditory inputs separately using dedicated encoders before fusing the features for contrastive learning during training. IS3 also includes an Intersection-over-Union (IoU) loss and a semantic alignment loss to improve localization accuracy during supervised training. The model is trained on the FlickrSoundNet aytar2016soundnet and VGG-Sound vggsound datasets.
IS3-FT is introduced to examine how training data quality and audio input format (stereo vs. mono) affect model behavior. We fine-tune the pre-trained IS3 is3 on our AudioCOCO training set (see Fig. 3). For the mono condition (IS3-FT-Mono), we average the two stereo channels into a single-channel input, which is then passed to the pre-trained audio encoder. For the stereo condition (IS3-FT-Stereo), we add a 1D convolutional layer before the audio encoder to fuse stereo features. In both setups, all network parameters are made learnable during fine-tuning.
Random is a chance model that randomly selects a location on the image as the predicted sound source location. It serves as a lower bound for SSL without using any audio-visual information.
@@
For both IS3 and IS3-FT models, feature maps from the final layers of the encoders are extracted, and cosine similarity is computed between them to produce a similarity heatmap. The predicted sound source location is identified as the point with the highest activation on the heatmap. See Supp Sec. S3 for how the sound source location is predicted by the transformer-based models.
@@
Evaluation metrics. To disentangle spatial localization from semantic alignment between visual and audio modalities, we define two metrics: Audio Accuracy (A-Acc) measures whether the model or human localizes the true sound source regardless of matching semantics. A-Acc = 1 if the peak activation falls within the segmentation mask of the sounding object; 0 otherwise. Vision Accuracy (V-Acc) measures alignment with visual semantics. V-Acc = 1 if the peak activation falls within any object of the correct category, even if it might not be the actual sound source, such as in MultiInstLoc conditions. In VOnly condition, V-Acc = 1 if the activation overlaps with any object from the 12 sound-emitting categories in AudioCOCO.
@@ 3 Results
We present the results of humans and AI models across all experimental conditions. Due to space constraints, we highlight comparisons between humans and the top performing AI models IS3 and IS3-FT-Stereo. Results for the remaining models are provided in Supp Tab. S1.
@@
Object size matters for humans and AI. As shown in Fig. 5, human A-Acc improves steadily with increasing object size, indicating that larger sounding objects are easier for humans to localize. Similarly, IS3-FT-Stereo shows a notable improvement in A-Acc from 19.3% to 52.2% as object size increases, though it still lags behind human performance. Among all models, IS3-FT-Stereo consistently outperforms the second-best model, IS3, across all object sizes. This is likely because IS3 is trained on datasets biased toward large, centered objects, whereas AudioCOCO offers greater diversity in sizes of sounding objects and their spatial locations. Despite these differences, both humans and all AI models achieve above-chance A-Acc when localizing large objects. However, only humans and IS3-FT-Stereo maintain consistently strong performance across all object sizes, including small objects, whereas IS3 does not.
@@
Figure 6: Humans and stereo-trained models show horizontal-vertical asymmetry in SSL accuracy, while mono-trained models do not. We report the proportion of trials (object size 2) where the predicted sound location falls within 6 degrees of visual angle from the ground truth, separately along the horizontal (textured bars) and vertical (plain bars) directions. See Supp Tab. S1 for results on object sizes 1 and 3.
@@
Figure 7: AI models show a bias towards objects that emit sounds, even in the absence of sound or in the presence of noisy sound. The V-Acc of AI models is presented in the VOnly condition, where either no sound or only noisy sound is present. Despite the lack of auditory cues, V-Acc of AI models on object size 2 remains higher than random (dark gray), indicating a bias towards audible objects.
@@
Conflict cues harm more than the lack of cues for AI. As shown in Fig.5, ConflictVCue and AbsVCue impair SSL performance for both humans and AI models, compared to the congruent conditions. While humans and IS3-FT-Stereo can still outperform chance, IS3 drops to near chance performance under these conditions. This suggests that humans and IS3-FT-Stereo are more robust to SSL with conflicting or missing visual cues. Interestingly, for IS3-FT-Stereo, conflicting visual cues cause a greater performance decline than the absence of visual cues that align with the sound source. This implies that IS3-FT-Stereo can be misled by conflicting visual cues. These findings suggest that, unlike humans, current AI models still place greater reliance on visual input for SSL.
AI fails when there are only auditory cues but humans can. In Supp Fig. S5, humans achieve A-Acc above chance even without any visual information (e.g., grayscale or Gaussian noise backgrounds), confirming that humans can perform SSL based on auditory cues alone. However, their performance is notably lower than in the congruent condition, highlighting that congruent visual information enhances SSL for humans. In contrast, IS3-FT-Stereo and IS3 performs worse than humans and drops to chance-level performance in AOnly condition, recapitulating that they depend heavily on congruent visual cue to support SSL.
In neuroscience francl2022deep , horizontal-vertical asymmetry in auditory precision describes the tendency for better localization in the horizontal than vertical axes. We measured the proportion of trials based on the Euclidean distance between predicted and ground-truth sound locations along each axis. As shown in Fig. 6, humans show a strong horizontal bias, with 52.1% of predictions falling within 6 degrees of visual angle horizontally, compared to 31.5% vertically. Remarkably, IS3-FT-Stereo—despite no exposure to human data—exhibits a similar, though weaker, bias. In contrast, IS3 shows no clear directional preference, likely reflecting its limited use of spatial stereo cues.
AI biases towards sound-emitting objects. Previous studies cavp ; is3 report that AI models often rely on visual shortcuts—favoring large, centered objects—due to dataset biases. To probe additional behavioral tendencies, we evaluate models under a vision-only condition. Without meaningful sound, any object could plausibly be the source. As shown in Fig. 7, models tend to favor sound-emitting objects (e.g., people, animals) over irrelevant ones (e.g., sky, trees), yielding significantly above-chance V-Acc. This suggests models carry prior knowledge about which objects typically produce sound. Interestingly, even noise input increases V-Acc compared to silence, indicating that audio noise may enhance this bias for AI. IS3-FT-Stereo exhibits the strongest effect, with substantially higher V-Acc than IS3.
@@
gao2024avsegformer
SSLTI
ssltie
FNAC
fnac
IS3-FT
-Stereo
Human    Acc(%) Rand
IS3
is3
IS3-FT
A- Acc Size1 1.6 4.2 5.9 5.5 0.6 0.7 8.9 29.1  Mono Size1 1.6 9.8 13.1
Size2 9.1 10.3 11.5 11.3 2.5 3.9 14.3 63.9  Size3 9.5 17.9 25.6
Size3 21.3 22.4 27.4 26.5 24.0 26.3 31.5 71.1  Size3 21.8 35.0 44.3
V- Acc Size1 8.4 17.2 20.5 19.6 13.2 13.4 85.4 97.5  Stereo Size1 1.6 11.5 17.5
Size2 17.8 22.8 26.0 24.7 20.7 25.5 89.5 99.1  Size2 9.5 21.3 30.0
Size3 26.2 49.5 59.5 56.3 49.4 52.8 93.7 99.8  Size3 21.9 42.0 45.7
@@
Table 1: Multi-instance SSL remains a challenging task for both humans and AI models. The table on the left summarizes the audio and visual localization accuracy of humans and AI models under the multi-instance condition across all object sizes. For AI models, both input data quality and the use of stereo audio substantially impact SSL performance. The table on the right reports the overall SSL performance in Audio Accuracy (A-Acc) of AI models trained with either mono or stereo audio described in Sec.2.2, excluding the VOnly condition. In both tables, bold and underlined values indicate the best and second-best performances, respectively.
@@
Multi-instance SSL remains challenging for humans and AI. As shown in Tab. 1a, V-Acc is high for both IS3-FT-Stereo and humans, indicating strong alignment between audio and visual semantics. This allows them to use audio cues to identify all visual objects with matching semantics. However, in scenes with multiple object instances, successful localization requires fine-grained SSL, beyond semantic matching. In these cases, A-Acc drops for both humans and IS3-FT-Stereo. Compared to the Congruent condition (with a single sounding object), multi-instance SSL is notably harder, as it demands precise spatial disambiguation. Despite this, both IS3-FT-Stereo and humans still perform above chance, especially for small objects. This demonstrates an ability to localize sounds at fine spatial resolution for both humans and IS3-FT-Stereo. Nonetheless, IS3-FT-Stereo still lags behind human performance, with the gap more pronounced in A-Acc than in V-Acc, underscoring the need for improved SSL in AI. Among all models, IS3-FT-Stereo performs best. It even outperforms strong baselines like CAVP and AVSeg, which are trained on large-scale, standard audio-visual datasets.
@@
Training data quality and stereo input are important for SSL. We report the average A-Acc across all object sizes for Random, IS3, and two fine-tuned variants of IS3 over five experimental conditions in Tab. 1b and for all other AI models in Supp Tab. S1 and Supp Sec. S4. IS3-FT-Stereo consistently outperforms all baselines by a large margin. Tab. 1 also shows that IS3-FT-Mono outperforms the original IS3 even with fewer examples to fine-tune on, highlighting the impact of training data quality. Moreover, IS3-FT-Stereo surpasses IS3-FT-Mono, underscoring the importance of stereo inputs in capturing spatial localization cues.
We present qualitative results in Fig. 8, comparing predicted sound source locations from humans, IS3, and IS3-FT-Stereo. IS3-FT-Stereo localizes sounds more accurately and often matches human predictions. In contrast, IS3 struggles with small or off-center objects. For example, IS3 fails to localize a dog in the top-right of an image in the Congruent condition and it frequently mislocalizes sources in incongruent conditions, such as horses, dogs, or cats. This suggests IS3 overly relies on visual cues and fails when auditory signals dominate. Nonetheless, IS3-FT-Stereo still underperforms humans in some cases. Additional qualitative examples are shown in Supp Fig. S6, indicating that IS3-FT-Stereo lacks the flexibility to resolve audio-visual conflicts or rely solely on audio when visual cues are absent.
@@
Figure 8: IS3 struggles to localize sound sources, whereas humans and IS3-FT-Stereo perform well across all four experimental conditions. The leftmost images (Columns 1 and 4) show the correct localization results made by human participants. Red boxes mark the ground truth sound source locations, while green circles indicate mouse click responses from multiple participants. The middle columns (2 and 5) and rightmost columns (3 and 6) display heatmaps predicted by IS3-FT-Stereo and IS3, respectively. Yellow triangles on the heatmaps denote the predicted sound source locations from each model. See the colorbar for the activation values of the heatmaps.

### v2

@@ 2.2 Experimental conditions in the AudioCOCO test set
Vision Only (VOnly) exploits multi-modal biases or priors. In this condition, only visual scenes are provided to the AI models, and they must localize the sounding object despite the absence of any meaningful sound. The audio could either be completely silent or filled with random Gaussian noise (mean 0, standard deviation 1). For example, the same visual stimulus as in the congruent condition is presented, but the dog sound is replaced with silence or noise (Fig. 2(e)).
Multi-Instance Localization (MultiInstLoc) follows the same motivation as the cocktail party problem haykin2005cocktail and features several objects of the same category, but the audio corresponds to only one specific instance, testing localization accuracy in a multi-instance scenario. For example, as illustrated in Fig. 2(f), a dog sound is played at the location of the left dog, while both dogs are visually present in the scene. This condition follows the same setup as the Congruent condition, but is more challenging due to the presence of multiple visually relevant objects. We selected images from the test set of the MSCOCO dataset, where the number of object instances within a given image is restricted to between 2 and 5 for the target category.
@@ 2.3 Human psychophysics experiment
We conducted in-lab psychophysics experiments on the AudioCOCO test set with 14 participants, collecting a total of 2,100 trials. All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution. Every experiment lasted approximately 40 minutes. The experimental setup is schematically illustrated in Fig. 2(h). Each trial began with a fixation cross displayed for 500 milliseconds, followed by the presentation of an image and audio pair drawn from one of the six experimental conditions (see Sec 2.2). The 6-second audio clip paired with the image stimulus continuously loops until the trial concludes. Participants were instructed to use a computer mouse to click on the perceived location of the sound source within a time limit of 20 seconds, while wearing stereo headphones throughout the experiment. All trials were randomly sampled, and their presentation sequences were shuffled to minimize order effects. If participants failed to respond within the allotted 20 seconds, the trial automatically ended, and the next trial commenced. Instead of relying on the pre-rendered image-audio pairs, we conducted an audio calibration procedure at the start of the experiment to account for individual differences in auditory perception. To achieve this, we implemented real-time stereo audio synthesis in Unity. Following calibration, an audio validation task was conducted to ensure successful calibration and adequate spatial hearing accuracy from the participants. See Supp Sec. S2 and Supp Fig.S6 for additional details on the human psychophysics experiments.
Figure 3: Overview of our neuroscience-inspired EchoPin model. EchoPin takes as input a static image paired with a two-channel stereo audio signal. The stereo waveforms are first filtered using the Head-Related Transfer Function (HRTF) to simulate sound filtering by the pinnae, and then converted into cochleagrams to mimic auditory processing in the human cochlea. The audio and visual streams are independently processed through dedicated encoders. During training, semantic alignment between the two modalities is enforced using a contrastive loss applied to paired audio and visual embeddings, while localization alignment is achieved by regressing the predicted sound source location (indicated by a red triangle) from the multimodal feature similarity map to the ground-truth location (red bounding box).
2.4 Our neuroscience-inspired EchoPin model
We introduce EchoPin, a neuroscience-inspired model for SSL (Fig. 3). The model takes image–audio pairs as input and emulates the human auditory periphery to decompose sounds in raw waveforms into frequency components. These representations are then aligned with visual features through a dual-encoder architecture for auditory and visual processing.
In human auditory neuroscience, incoming sound waveforms are directionally shaped by the pinna, head, and torso. These spectral transformations encode elevation-specific notches and interaural differences, which are essential for resolving front–back and vertical ambiguities van2019cortical . For implementation, we use pre-measured Head-Related Transfer Functions (HRTFs) from human ears, developed in Unity simulations and based on the extensive KEMAR dummy head dataset gardner1994hrft . KEMAR is equipped with microphones in the ear canals to capture how sounds from different directions are filtered by the head and pinnae. The dataset includes left and right ear impulse responses recorded from a Realistic Optimus Pro 7 loudspeaker positioned 1.4 meters from KEMAR, covering 710 spatial positions with elevations from $40^{\circ}$ to $90^{\circ}$.
Next, the HRTF-filtered time-domain sound waveform is passed through a cochlear-inspired frequency decomposition, converting it into a cochleagram using the PyCochleagram library gardner1994hrft . The cochleagrams are constructed via Equivalent Rectangular Bandwidth (ERB) filterbanks, capturing the tonotopic and temporal resolution of sound across frequency channels. This representation retains key auditory features, including pitch, timbre, and spatial cues. The resulting 10-second stereo waveform, sampled at 16 kHz, is transformed into cochleagrams, yielding a tensor of size $66\times 160{,}000\times 2$, where the dimensions correspond to 66 ERB filters (after truncating 10 high-frequency channels for efficiency), 160k temporal samples, and two binaural channels. The binaural channels are first integrated using 1D convolution kernels to merge information across ears and then fed into the dual-encoder architecture of IS3 (is3, ), allowing separate audio and visual processing streams.
During training, we initialize all weights from the pre-trained IS3 model, except for the first 1D-convolution layer in the audio encoder described above, and then optimize all parameters end-to-end using supervised learning. Two losses in the IS3 (is3, ) model are employed: (i) a Triplet loss to enforce semantic alignment by pulling matched audio–visual embeddings closer than mismatched ones, and (ii) a CIoU loss to penalize spatial deviation between predicted and ground-truth sounding-object bounding boxes. This combination enables EchoPin to jointly capture what is sounding and where it is located. See Supp Sec. S3 for extra implementation details.
Model variants of EchoPin. To study the effects of design components in EchoPin, we introduce two model variants: EchoPin-M (Mono) averages the two HRTF-filtered stereo channels into a single time-domain waveform, allowing us to examine the impact of mono versus stereo audio on SSL tasks. EchoPin-S (Stereo) uses the HRTF-filtered stereo waveforms as input, but processes them with standard mel-spectrograms instead of cochleagrams. Comparing these variants with the full EchoPin model allows us to examine how stereo structure captured by the pinnae and frequency-specific cochlear modulation affect SSL performance. See Tab. 1(b) and Sec. 3 for results and discussion.
2.5 Baseline methods and evaluation metrics
We benchmark EchoPin and the state-of-the-art multimodal models, including SSLTI (ssltie, ), LVS (vggss, ), FNAC (fnac, ), CAVP (cavp, ), AVSegformer (gao2024avsegformer, ), IS3 (is3, ), ImageBind (imagebind, ), and LanguageBind (languagebind, ), using the same stimuli as in our human psychophysics experiments. While humans can leverage the pinna, head, and torso to encode elevation-specific auditory cues, models lack these physical structures. To ensure fair comparisons across models and with human participants, all AudioCOCO audios are HRTF-filtered to approximate human auditory processing, and these filtered sounds are used for all model evaluations. In the main text, we provide brief overviews of IS3 (is3, ) and a random baseline, and report their performance alongside our proposed EchoPin model. Detailed descriptions of the other models and their extended experimental results are provided in Supp. Sec. S3.
@@ 2.5 Baseline methods and evaluation metrics
IS3 is3 is a dual-stream architecture with 2D CNNs, which processes visual and monaural auditory inputs separately using dedicated encoders before fusing the features for contrastive learning during training. IS3 also includes an Intersection-over-Union (IoU) loss and a semantic alignment loss to improve localization accuracy during supervised training. The model is trained on the FlickrSoundNet aytar2016soundnet and VGG-Sound vggsound datasets. Random is a chance model that randomly selects a location on the image as the predicted sound source location. It serves as a lower bound for SSL without using any audio-visual information.
@@
Predicting target sound locations. For IS3, EchoPin, and other 2D CNN–based models, feature maps from the final layers of the visual and audio encoders are extracted, and cosine similarity is computed between them to generate a similarity heatmap. The predicted sound location is taken as the point with the highest activation on this heatmap. See Supp. Fig. S7 for an illustration of how transformer-based baselines predict target sound source locations.
@@
Evaluation metrics. To disentangle spatial localization from semantic alignment between visual and audio modalities, we define two metrics: Audio Accuracy (A-Acc) measures whether the model or human localizes the true sound source regardless of matching semantics. A-Acc = 1 if the peak activation falls within the bounding box of the sounding object; 0 otherwise. Vision Accuracy (V-Acc) measures alignment with visual semantics. V-Acc = 1 if the peak activation falls within any object of the correct category, even if it might not be the actual sound source, such as in MultiInstLoc conditions. In VOnly condition, V-Acc = 1 if the activation overlaps with any object from the 12 sound-emitting categories in AudioCOCO.
To robustly evaluate model performance, we consider three complementary factors that could influence results in Supp Sec. S5. First, A-Acc can be biased by object size, so we introduce a chance-corrected A-Acc (Supp Tab. S5; Supp Tab. S1). Second, human clicks may fall near but not inside the target, so we treat clicks within a thresholded radius as correct in Supp. Tab. S2. None of these metric variants alter the conclusions. Finally, we evaluate the predicted target object bounding boxes by all the models using corrected Intersection over Union (cIoU, zheng2021enhancing ), where EchoPin continues to outperform other baselines (Supp Sec. S5 and Supp Tab. S3).
@@ 3 Results
We report results from both human participants and AI models across all experimental conditions and object sizes. For brevity, the main text focuses on comparisons between humans and the two best-performing models, IS3 and EchoPin. Additional qualitative analyses for all other models are presented in Supp. Tab. S4.
@@
Figure 6: EchoPin shows human-like horizontal-vertical asymmetry in SSL accuracy, while other models do not. We report the proportion of trials (object size 2) where the predicted sound location falls within 6 degrees of visual angle from the ground truth, separately along the horizontal (textured bars) and vertical (plain bars) directions.
@@
Figure 7: AI models show a bias towards objects that emit sounds, even in the absence of sound or in the presence of noisy sound. The V-Acc of AI models is presented in the VOnly condition, where either no sound or only noisy sound is present. Despite the lack of auditory cues, V-Acc of AI models on object size 2 remains higher than random (dark gray).
Object size matters for humans and AI. As illustrated in Fig. 5, human A-Acc steadily increases with object size, suggesting that larger sounding objects are easier for humans to localize. EchoPin exhibits a similar trend, with A-Acc rising substantially from 11.2% to 76.0% as object size grows. Across all object sizes, EchoPin consistently outperforms IS3 and the other models. This difference likely arises because IS3 is trained on datasets biased toward large, centered objects, whereas AudioCOCO provides a wider range of object sizes and spatial locations. While both humans and AI models perform above chance for large objects, only humans and EchoPin maintain strong, consistent performance across all object sizes, including smaller ones—performance that IS3 fails to achieve.
@@
Conflict cues harm more than the lack of cues for AI. As shown in Fig. 5, both the ConflictVCue and AbsVCue conditions impair SSL performance in humans compared to the congruent conditions. While humans and EchoPin still perform significantly above chance, IS3 drops to near-chance levels under the ConflictVCue condition. This suggests that humans and EchoPin are more robust in SSL tasks involving conflicting or missing visual cues, whereas IS3 relies heavily on visual information. Notably, unlike humans, EchoPin shows a greater performance decline when visual cues are conflicting but not when they are absent, indicating that it is more easily misled by incongruent visual information yet remains stable in the absence of such cues.
@@
EchoPin Human    Acc(%) Rand IS3 EchoPin-M/S EchoPin
A- Acc Size1 1.6 4.8 2.9 2.5 4.5 25.7  Mono Size1 1.6 3.0 3.6 -
Size2 9.1 7.9 7.5 7.3 24.1 36.4  Size2 9.4 13.9 15.8 -
Size3 21.3 22.4 20.4 20.2 47.1 38.6  Size3 19.8 28.7 31.4 -
V- Acc Size1 8.4 11.9 10.5 11.2 37.5 60.9  Stereo Size1 1.6 - 5.3 9.7
Size2 17.8 24.1 23.0 23.7 53.8 82.8  Size2 9.4 - 17.0 31.3
Size3 26.2 40.9 39.5 40.9 64.2 89.1  Size3 19.8 - 35.2 47.6
@@
Table 1: Multi-instance SSL remains a challenging task for both humans and AI models. The table (a) on the left summarizes the audio and visual localization accuracy of humans and AI models under the multi-instance condition across all object sizes. For AI models, both input data quality and the use of stereo audio substantially impact SSL performance. Table (b) on the right summarizes the average A-Acc across the Congruent, ConflictVcue, AbsVcue, and AOnly conditions for models trained with either mono or stereo audio. The second-to-last column shows the results of EchoPin-M (Rows 1–3) and EchoPin-S (Rows 4–6). See Sec. 2.4 for details on these variants. (–) indicates that the results are not applicable due to the model configurations. In both tables, best is in bold and the second best is underlined.
AI fails when there are only auditory cues but humans can. As shown in Supp. Fig. S8, humans achieve above-chance A-Acc even without visual information (e.g., gray or Gaussian noise backgrounds), confirming that they can perform SSL based solely on auditory cues. However, their performance remains lower than in the congruent condition, indicating that congruent visual information facilitates SSL. Similar trends are observed in IS3 and EchoPin, with EchoPin consistently outperforming IS3 under both AOnly conditions (gray or Gaussian background). Interestingly, both humans and EchoPin perform better with gray than Gaussian backgrounds, suggesting that incongruent visual noise can distract attention and interfere with SSL, whereas a neutral (gray) background minimizes interference.
EchoPin shows human-like asymmetry in auditory spatial precision. In neuroscience, auditory spatial precision is known to exhibit a horizontal–vertical asymmetry: humans localize sounds more accurately along the azimuth (horizontal) than the elevation (vertical) axis francl2022deep . To quantify this effect, we measured the proportion of trials where predictions fell within six degrees of visual angle from ground-truth locations, separately for horizontal and vertical dimensions (Fig. 6). As expected, humans showed a strong horizontal advantage, localizing targets in 86.1% of trials horizontally but only 66.7% vertically. Remarkably, EchoPin exhibited a similar asymmetry pattern despite being trained without any human behavioral data. In contrast, IS3, which relies solely on monaural Mel-spectrograms, also showed asymmetry but to a much lesser degree. Although both models have benefited from spatial filtering by the pinnae, the observed asymmetry in EchoPin arises from its biologically grounded auditory frequency decomposition in the cochlea and its stereo audio perception. To further validate this, we introduced an EchoPin variant, EchoPin-Ro, in which the interaural axis was rotated by 90 degrees in Unity, effectively simulating vertically aligned ears. When audio was re-rendered under this configuration, the model’s asymmetry was reversed, confirming the structural origin of this effect.
AI biases toward sound-emitting objects. Previous studies cavp ; is3 show that AI models often exploit visual shortcuts by favoring large or centered objects. To examine additional behavioral biases, we evaluate models under the VOnly condition. Without meaningful sound, any object could plausibly be the sound source. As shown in Fig. 7, models tend to localize sounds to vocal objects (e.g., people, animals) rather than irrelevant ones (e.g., sky, trees), resulting in above-chance V-Acc. This indicates that models encode prior knowledge of which objects typically produce sound. Furthermore, for all the models, Gaussian noise audio input leads to lower V-Acc than the absence of audio, indicating that even noisy audio can reduce the models’ over-reliance on visual signals.
@@
Multi-instance SSL remains challenging for humans and AI. As shown in Tab. 1a, V-Acc is high for both EchoPin and humans, indicating strong alignment between audio and visual semantics. This allows them to use audio cues to identify all visual objects with matching semantics. However, in scenes with multiple object instances, successful localization requires fine-grained SSL, beyond semantic matching. In these cases, A-Acc drops for both humans and EchoPin. Compared to the Congruent condition (with a single sounding object), multi-instance SSL is notably harder, as it demands precise spatial disambiguation. Despite this, humans still perform above chance, especially for small objects. Similarly, EchoPin achieves high A-Acc on object sizes 2 and 3. However, it still lags behind human performance, with the gap more pronounced in small object size. Despite this, both EchoPin and humans still perform above chance, especially for small objects. This demonstrates an ability to localize sounds at fine spatial resolution for both humans and EchoPin. Among all models, EchoPin performs best. It even outperforms strong baselines like CAVP and AVSeg, which are trained on large-scale, standard audio-visual datasets.
@@
Training data quality and stereo input are important for SSL. We report the average A-Acc across all object sizes for Random, IS3, and the EchoPin variants under four experimental conditions in Tab. 1b. EchoPin consistently achieves the highest performance across all conditions. Notably, EchoPin-M surpasses IS3 despite using fewer fine-tuning examples, underscoring the importance of high-quality training data. Moreover, EchoPin-S further improves over EchoPin-M, highlighting the advantage of incorporating human-like stereo configurations for spatial localization.
We further visualize the predicted sound source locations from human participants, IS3, and EchoPin in Fig. 8. EchoPin localizes sound sources more accurately and often aligns closely with human judgments. In contrast, IS3 struggles with small or peripheral targets—for instance, it fails to localize a dog in the top-left corner under the Congruent condition and frequently misattributes sounds to other vocal objects (e.g., person, motorbike). Although EchoPin markedly improves SSL robustness, it is still inferior to human performance in complex scenes. Failure cases and additional comparisons with other baselines, such as CAVP, are provided in Supp. Sec. S4, Supp. Fig. S9, and Supp. Fig. S10.
@@
Figure 8: IS3 struggles to localize sound sources, whereas humans and EchoPin perform well across all four experimental conditions. The leftmost images (Columns 1 and 4) show the correct localization results made by human participants. Red boxes mark the ground truth sound source locations, while green circles indicate mouse click responses from multiple participants. The middle columns (2 and 5) and rightmost columns (3 and 6) display heatmaps predicted by EchoPin and IS3, respectively. Yellow triangles on the heatmaps denote the predicted sound source locations from each model. See the colorbar for the activation values of the heatmaps.